ebeebe:

Write the standard form of the line that contains a y-intercept of -3 and passes through the point (3, 0). Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution.

Enorym:

Welcome to QC The slope-intercept equation of a line is \(y=mx+b\) m = slope b = y-intercept The question gave us the y-intercept (0,-3), so we can substitute that for b \(y=mx-3\) They also gave us a point (3,0) [Remember that points are written as (x,y) which is also included in the equation] Substitute this into the equation we have \(0=m(3)-3\) => \(0=3m-3\) Now you need to simplify by isolating the m If you are unsure on how to do this feel free to ask